OpenStudy (anonymous):

how do you go about simplifying -7√ 216x

OpenStudy (anonymous):

\[216=3*3*2*2*6=6\sqrt{6}\]

OpenStudy (anonymous):

I'll let you solve the rest, ask me if something is not clear

OpenStudy (anonymous):

i got -21√ 24x am i right??

OpenStudy (anonymous):

is x inside of the square root

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

is that wrong

OpenStudy (anonymous):

You multiply stuffs who are outside of square root together, and multiply stuffs who are inside of the square root together. It should give you -42 * sqrt(6x)

OpenStudy (anonymous):

oooh okay i see what i did thanks so much

OpenStudy (anonymous):

Np